OpenStudy (anonymous):

A hungry person eat 15 hot dogs in 4 minutes. How long will it take to eat 25 hot dogs? DO SOVE THE PROBLEM, JUST SET UP THE PROPORTION

OpenStudy (anonymous):

15/4 = 25/x

OpenStudy (anonymous):

15x=100

OpenStudy (anonymous):

x=100/15

OpenStudy (anonymous):

x=6.666

OpenStudy (anonymous):

so I would divide?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Im not sure how you did this but thanks

OpenStudy (anonymous):

where did the 100 come from?

OpenStudy (anonymous):

?